# Scoring Nim

### v1

@@ 5 General Properties of the Payoff Function
When there are three or more piles, explicitly determining $f_{N}(\bm{p})$ becomes challenging. In this section, we establish several propositions regarding $f_{N}(\bm{p})$ in Scoring Nim. The following two propositions are obtained by dividing the piles into two groups and analyzing the game as a disjunctive sum.
Proposition 5.1.
@@
Proof. Let $\bm{p}=(p_{1},p_{2},\ldots,p_{n},1,1)$. Divide the piles into $\bm{\overline{p}}=(p_{1},p_{2},\ldots,p_{n})$ and $\bm{\underline{p}}=(1,1)$, and let $\alpha=f_{N}(\bm{\overline{p}})$.
We first show that the second player can ensure $f_{N}(\bm{p})\leq\alpha$. If the first player makes a move in $\bm{\overline{p}}$, the second player responds optimally within $\bm{\overline{p}}$. If the first player makes a move in $\bm{\underline{p}}$, the second player removes the remaining stone from $\bm{\underline{p}}$. Furthermore, if the first player takes the last stone from $\bm{\overline{p}}$, the second player removes a remaining stone from $\bm{\underline{p}}$, if possible. In this case, the first player will then remove the remaining stone from $\bm{\underline{p}}$, ending the game. As a result, the stones in $\bm{\underline{p}}$ are removed one by one by each player, meaning that the identity of the player who takes the last stone in $\bm{\overline{p}}$ remains the same in $\bm{p}$. Therefore, the second player can ensure that $f_{N}(\bm{p})\leq\alpha$.
Similarly, we show that the first player can ensure $f_{N}(\bm{p})\geq\alpha$. The first player begins by making the optimal move in $\bm{\overline{p}}$. If the second player makes a move in $\bm{\overline{p}}$, the first player responds optimally within $\bm{\overline{p}}$ if there are stones remaining in $\bm{\overline{p}}$. If the second player makes a move in $\bm{\underline{p}}$, the first player removes the remaining stone from $\bm{\underline{p}}$. By following this strategy, the first player ensures $f_{N}(\bm{p})\geq\alpha$.  $\Box$
The following proposition states that if a position is symmetric, meaning it consists of two identical groups of piles, the same payoff function as (3) is obtained.
Proposition 5.2.
Let the number of piles $n$ be even (i.e., $n=2k$), and suppose that in $\bm{p}=(p_{1},p_{2},\ldots,p_{n})$, the condition $p_{2j-1}=p_{2j}$ holds for $1\leq j\leq k$. If there exists $i$ such that $p_{i}\geq 2$,
$f_{N}(\bm{p})=1-|1+N|$
@@
Proof. By Proposition 5.1, we may assume that all $p_{i}$ are at least 2. For $k=1$, the result follows directly from (3). We prove the claim for $k\geq 2$ by induction on $|\bm{p}|$. Let $\alpha=1-|1+N|$.
First, we establish a strategy for the second player to ensure $f_{N}(\bm{p})\leq\alpha$. When the first player takes stones from a pile containing $p_{i}$ stones, the second player responds by taking the same number of stones from another pile that also originally contained $p_{i}$ stones. This restores the symmetry of the position, and by the induction hypothesis, we conclude that $f_{N}(\bm{p})\leq\alpha$.
Next, we construct a strategy for the first player to ensure $f_{N}(\bm{p})\geq\alpha$. The first player begins by taking all stones from the pile with the largest number of stones. If the second player then takes the same number of stones from another pile, symmetry is restored. If instead, the second player takes fewer stones than the first player on the initial move, the first player can continue to always take all stones from the pile with the largest remaining number of stones. By the end of the game, the total number of stones taken by the first player will be at least two more than that taken by the second player, ensuring that the final score of the first player is at least $2-|N|$, which is at least $\alpha$.  $\Box$
@@
Proposition 5.3.
Let $\bm{p}=(p_{1},p_{2},\ldots,p_{n})$ with $p_{1}\geq p_{2}\geq\cdots\geq p_{n}>0$. Then, for $-1\leq N\leq 0$, the following hold:
(i) When $n$ is odd, i.e., $n=2k-1$, we have
$f_{N}(\bm{p})=\sum_{j=1}^{k-1}\left(p_{2j-1}-p_{2j}\right)+p_{2k-1}+N.$
@@
$f_{N}(\bm{p})=\sum_{j=1}^{k}\left(p_{2j-1}-p_{2j}\right)-N.$
Proof. Scoring Nim for $-1\leq N\leq 0$ can be viewed as a variant of the stone-taking game in which only the last stone is worth $1+N$ points (with $0\leq 1+N\leq 1$) instead of 1. Since each player aims to take as many stones as possible, this proposition follows almost trivially.  $\Box$
@@
Definition 5.4 (P-position set in Nim).
Define the sets of positions $\bm{p}=(p_{1},p_{2},\ldots,p_{n})$ as follows:
$P^{+}$ $=$ $\{p_{1}\oplus p_{2}\oplus\cdots\oplus p_{n}=0\},$
$P^{-}$ $=$ $\{\mbox{there exists }i\mbox{ such that }p_{i}\geq{2},\mbox{ and }p_{1}\oplus p_{2}\oplus\cdots\oplus p_{n}=0\}$
$\cup\ \{p_{i}\leq 1\mbox{ for all }i,\mbox{ and }|\bm{p}|\mbox{ is odd}\}.$
@@
Proposition 5.5.
(i) There exists an integer $c^{+}(\bm{p})$ such that
$f_{N}(\bm{p})=\left\{\begin{array}[]{ll}c^{+}(\bm{p})-N&\mbox{ if }\bm{p}\in P^{+},\\ c^{+}(\bm{p})+N&\mbox{ if }\bm{p}\not\in P^{+}\end{array}\right.$
holds for $N\geq|\bm{p}|-2$.
(ii) There exists an integer $c^{-}(\bm{p})$ such that
$f_{N}(\bm{p})=\left\{\begin{array}[]{ll}c^{-}(\bm{p})+N&\mbox{ if }\bm{p}\in P^{-},\\ c^{-}(\bm{p})-N&\mbox{ if }\bm{p}\not\in P^{-}\end{array}\right.$
holds for $N\leq-|\bm{p}|$.
@@
This proposition states that, for example, when $N$ is positive and sufficiently large, the player who takes the last stone is determined by whether the position is in $P^{+}$. Under the constraint that the player taking the last stone must always move to a P-position, the difference in the number of stones taken by both players assuming they take as many stones as possible is given by $c^{+}(\bm{p})$. For instance, in the position $(5,4,2)$, if $N$ is positive and sufficiently large, the first player will take the last stone. However, to achieve this, the first player must allow the second player to take 7 stones, which is unavoidable (while first player takes 4 stones). Thus, the payoff function is given by $f_{N}(5,4,2)=(4-7)+N=-3+N$, implying that $c^{+}(5,4,2)=-3$.
@@
Proof of Proposition 5.5. (i) When $\bm{p}\not\in P^{+}$, the first player has a winning strategy in the normal play Nim (i.e., taking the last stone), ensuring they obtain at least $1+N$ points. On the other hand, if the first player does not take the last stone, the total points obtained by taking all remaining stones is at most $|\bm{p}|-1$. Thus, if $11+N\geq|\bm{p}|-$ 1, which simplifies to $N\geq|\bm{p}|-2$, the first player will always select to take the last stone. Under the constraint that the first player must always select the next game state from within $P^{+}$, an optimal strategy that maximizes their payoff is determined. The value of $c^{+}(\bm{p})$ is given by the difference between the number of stones the first player can take and the number the second player can take. If $\bm{p}\not\in P^{+}$, then $c^{+}(\bm{p})$ is determined under the constraint that the second player must always select the next game state from within $P^{+}$.
@@
(ii) In this case, similar to (i), we compare the maximum possible points when a player takes the last stone, $|\bm{p}|+N$, with the minimum possible points when a player avoids to take it, which is 0. If $|\bm{p}|+N\leq 0$, the player who wins in the misère play Nim will always select the next game state from within $P^{-}$. Under this constraint, $c^{-}(\bm{p})$ is determined.  $\Box$
@@
As an example of $c^{+}(\bm{p})$ and $c^{-}(\bm{p})$, we compute $c^{+}(3,2,1)$ and $c^{-}(3,2,1)$ in the three-pile Scoring Nim. Starting from $(3,2,1)$, the opponent responds with a move that leads to the next game state in either $P^{+}$ or $P^{-}$ position when $|N|$ is sufficiently large. Specifically, possible moves include $(3,2,1)\to(2,2,1)\to(2,2,0)\in P^{+}\cap P^{-}$, $(3,2,1)\to(1,2,1)\to(1,0,1)\in P^{+}$ or $(1,1,1)\in P^{-}$, among others. By evaluating all possible moves, we find that:
$c^{+}(3,2,1)=3-1+c^{+}(0,1,1)=2,\quad c^{-}(3,2,1)=3-2+c^{+}(0,0,1)=2.$
@@
Similarly, considering all possible outcomes for $(2k+1,2k,1)$ ($k\geq 2$) when $|N|$ is sufficiently large, we find that the optimal sequence of moves consists of the current player first taking 3 stones from the pile with $12k+$ 1 stones, followed by the opponent taking 1 stone from the pile with $2k$ stones. This leads to the inductive conclusion that:
$c^{+}(2k+1,2k,1)=c^{-}(2k+1,2k,1)=2k\quad(k\geq 1).$ (5)
This result will be used in the proof of Theorem 6.3 in the next section.
@@ 6 Changes in the Optimal Strategy Depending on the Bonus Value
Proof. For (6), we apply the recursion formula (2) and use induction on the total number of stones to establish the result. We assume here $x>z$.
First, the player may take all the stones from the pile containing $z$ stones. By (3), we obtain
@@
To establish the reverse inequality, we show that for any possible move by the first player, the opponent can always respond in such a way that $f_{N}(x,x,z)$ does not exceed $z-1+|1+N|.$
  • If the first player moves to $(x,0,z)$, then by (4), it follows that
    $x-f_{N}(x,0,z)=x-\left((x-z)+\bigl{|}1-|1+N|\bigr{|}\right)\leq z-(1-|1+N|).$
  • If the first player moves to $(x,x^{\prime},z)$ (where $0<x^{\prime}<x$), the opponent will respond with a move to $(x^{\prime},x^{\prime},z)$ except when $x^{\prime}=1$ and $z=1$. By the induction hypothesis, it follows that
@@
    If $x^{\prime}=1$ and $z=1$, the opponent will respond with a move to $(0,1,1)$, then it follows that
@@
  • If the first player moves to $(x,x,z^{\prime})$ (where $0<z^{\prime}<z$), the opponent will respond with a move to $(x,x,0)$. Then, by (3), it follows that
@@
Thus, the case $x>z$ has been proved. The cases $x=z$ and $x<z$ can be established in a similar manner.  $\Box$
@@
We classify $f_{N}(x,y,1)$ according to cases, where the condition $(x,y,1)\in P^{+}\cap P^{-}$ holds only for $(2k+1,2k,1)$ with $k\geq 1$. We define the function $F_{k}(N)$ for $k\geq 1$ as follows.
Definition 6.2.
@@
For example, when $k=3$, $F_{3}(N)$ is depicted in Figure 3. The function $F_{k}(N)$ attains a value of 0 when $N=\pm 2k$.
[…]
@@
Thus, we have fully characterized $f_{N}(x,y,1)$. For example, $f_{N}(8,7,1)$ is an instance of the case where $k=3$ and $x=8$ in part (iii) of the theorem above. It is given by $\max\{F_{3}(N),-4+|N|\}$, which is represented by the bold line in Figure 4.
[…]
Figure 4: $f_{N}(8,7,1)=\max\{F_{3}(N),-4+|N|\}$ (bold line)
One key observation concerns the number of breakpoints in $f_{N}(x,y,1)$, namely, points where the slope changes sign. We find that $f_{N}(2k+1,2k,1)$ has $34k-$ 3 breakpoints, indicating that the number of breakpoints can grow arbitrarily large as the number of stones increases.
Another important property concerns the maximum and minimum values of $f_{N}(x,y,1)$ with respect to $N$.
Corollary 6.4.
(i) For all $N$, it follows that $f_{N}(2k+1,2k,1)\leq 2$, with equality holding for all $N$ in $J_{k}$. Furthermore, if $N$ is an odd integer, then $f_{N}(2k+1,2k,1)\leq 1$, and if $|N|\geq 2k+1$, then $f_{N}(2k+1,2k,1)\leq-1$.
(ii) For $x\geq 2k+2$, it follows that $f_{N}(x,2k,1)\geq x-2k-1$ for all $N$, with equality holding only when $N=\pm 2k$.
(iii) For $x\geq 2k+2$, it follows that $f_{N}(x,2k+1,1)\geq x-2k-1$ for all $N$.
These results are used repeatedly in the proof of Theorem 6.3. Notably, in (ii), equality holds for only two values of $N$. Specifically, $f_{N}(6,4,1)$ is an instance of the case where $k=2$ and $x=6$ in Theorem 6.3(ii), and as illustrated in Figure 5, it attains its minimum value at $N=\pm 4$.
[…]
@@
Proof of Theorem 6.3. For $k=1$, (i), (ii), and (iii) can be verified directly. For $k\geq 2$, we proceed by induction, assuming that (i), (ii), and (iii) hold for all $k<m$, and proving that (i) for $k=m$.
As stated in Section 3, to establish (i), it suffices to consider integer values of $N$. We outline the proof strategy as follows:
  • By showing that $f_{N}(2m+1,2m,1)\geq 2$ when $N$ is an even integer in the range $-(2m-2)\leq N\leq 2m-2$, it follows that $f_{N}(2m+1,2m,1)\geq F_{m}(N)$ for all $N$.
  • By showing that $f_{N}(2m+1,2m,1)\leq 1$ when $N$ is an odd integer in the range $-(2m-3)\leq N\leq 2m-3$, and using (5), it follows that $f_{N}(2m+1,2m,1)\leq F_{m}(N)$ for all $N$.
First, when $N=\pm 2l$ and $0\leq l<m$, the first player has the option to move from the position $(2m+1,2m,1)$ to $(2l,2m,1)$. By the induction hypothesis, i.e., using Corollary 6.4(ii), we have $f_{\pm 2l}(2l,2m,1)=2m-2l-1$ so we obtain

### v2

@@ 5 General properties of the payoff function
When there are three or more piles, explicitly determining $f_{N}(\boldsymbol{p})$ becomes challenging. In this section, we establish several propositions regarding $f_{N}(\boldsymbol{p})$ in Scoring Nim. The following two propositions are obtained by dividing the piles into two groups and analyzing the game as a disjunctive sum.
Proposition 5.
@@
Proof. Let $\boldsymbol{p}=(p_{1},p_{2},\ldots,p_{n},1,1)$. Divide the piles into $\boldsymbol{\overline{p}}=(p_{1},p_{2},\ldots,p_{n})$ and $\boldsymbol{\underline{p}}=(1,1)$, and let $\alpha=f_{N}(\boldsymbol{\overline{p}})$.
We first show that the second player can ensure $f_{N}(\boldsymbol{p})\leq\alpha$. If the first player makes a move in $\boldsymbol{\overline{p}}$, the second player responds optimally within $\boldsymbol{\overline{p}}$. If the first player makes a move in $\boldsymbol{\underline{p}}$, the second player removes the remaining stone from $\boldsymbol{\underline{p}}$. Furthermore, if the first player takes the last stone from $\boldsymbol{\overline{p}}$, the second player removes a remaining stone from $\boldsymbol{\underline{p}}$, if possible. In this case, the first player will then remove the remaining stone from $\boldsymbol{\underline{p}}$, ending the game. As a result, the stones in $\boldsymbol{\underline{p}}$ are removed one by one by each player, meaning that the identity of the player who takes the last stone in $\boldsymbol{\overline{p}}$ remains the same in $\boldsymbol{p}$. Therefore, the second player can ensure that $f_{N}(\boldsymbol{p})\leq\alpha$.
Similarly, we show that the first player can ensure $f_{N}(\boldsymbol{p})\geq\alpha$. The first player begins by making the optimal move in $\boldsymbol{\overline{p}}$. If the second player makes a move in $\boldsymbol{\overline{p}}$, the first player responds optimally within $\boldsymbol{\overline{p}}$ if there are stones remaining in $\boldsymbol{\overline{p}}$. If the second player makes a move in $\boldsymbol{\underline{p}}$, the first player removes the remaining stone from $\boldsymbol{\underline{p}}$. By following this strategy, the first player ensures $f_{N}(\boldsymbol{p})\geq\alpha$.  $\Box$
The following proposition states that if a position is symmetric, meaning it consists of two identical groups of piles, the same payoff function as (4.3) is obtained.
Proposition 6.
Let the number of piles $n$ be even (i.e., $n=2m$), and suppose that in $\boldsymbol{p}=(p_{1},p_{2},\ldots,p_{n})$, the condition $p_{2j-1}=p_{2j}$ holds for $1\leq j\leq m$. If there exists $i$ such that $p_{i}\geq 2$,
$f_{N}(\boldsymbol{p})=1-|1+N|$
@@
Proof. By Proposition 5, we may assume that all $p_{i}$ are at least 2. For $m=1$, the result follows directly from (4.3). We prove the claim for $m\geq 2$ by induction on $|\boldsymbol{p}|$. Let $\alpha=1-|1+N|$.
First, we establish a copy-cat strategy for the second player to ensure $f_{N}(\boldsymbol{p})\leq\alpha$. When the first player takes stones from a pile containing $p_{i}$ stones, the second player responds by taking the same number of stones from another pile that also originally contained $p_{i}$ stones. This restores the symmetry of the position, and by the induction hypothesis, we conclude that $f_{N}(\boldsymbol{p})\leq\alpha$.
Next, we construct a strategy for the first player to ensure $f_{N}(\boldsymbol{p})\geq\alpha$. The first player begins by taking all stones from the pile with the largest number of stones. If the second player then takes the same number of stones from another pile, symmetry is restored. If instead, the second player takes fewer stones than the first player on the initial move, the first player can continue to always take all stones from the pile with the largest remaining number of stones. By the end of the game, the total number of stones taken by the first player will be at least two more than that taken by the second player, ensuring that the final score of the first player is at least 2|N|2-|N|, which is at least $\alpha$.  $\Box$
@@
Proposition 7.
Let $\boldsymbol{p}=(p_{1},p_{2},\ldots,p_{n})$ with $p_{1}\geq p_{2}\geq\cdots\geq p_{n}>0$. Then, for $1\leq N\leq 0$, the following hold:
(i) When $n$ is odd, i.e., n=2k1n=2k-1, we have
$f_{N}(\boldsymbol{p})=\sum_{j=1}^{k-1}\left(p_{2j-1}-p_{2j}\right)+p_{2k-1}+N.$
@@
$f_{N}(\boldsymbol{p})=\sum_{j=1}^{k}\left(p_{2j-1}-p_{2j}\right)-N.$
Proof. Scoring Nim for $1\leq N\leq 0$ can be viewed as a variant of the stone-taking game in which only the last stone is worth $1+N$ points (with $0\leq 1+N\leq 1$) instead of 1. Since each player aims to take as many stones as possible, this proposition follows almost trivially.  $\Box$
@@
Definition 2 (P-position set in Nim).
Define the sets of positions $\boldsymbol{p}=$
$(p_{1},p_{2},\ldots,p_{n})$ as follows:
$P^{+}$ $=$ $\{\ \boldsymbol{p}\ |\ p_{1}\oplus p_{2}\oplus\cdots\oplus p_{n}=0\ \},$
$P^{-}$ $=$ $\{\ \boldsymbol{p}\ |\ \mbox{there exists }i\mbox{ such that }p_{i}\geq{2},\mbox{ and }p_{1}\oplus p_{2}\oplus\cdots\oplus p_{n}=0\ \}$
$\cup\ \{\ \boldsymbol{p}\ |\ p_{i}\leq 1\mbox{ for all }i,\mbox{ and }|\boldsymbol{p}|\mbox{ is odd}\ \}.$
@@
Proposition 8.
(i) There exists an integer $c^{+}(\boldsymbol{p})$ such that
$f_{N}(\boldsymbol{p})=\left\{\begin{array}[]{ll}c^{+}(\boldsymbol{p})-N&\mbox{ if }\boldsymbol{p}\in P^{+},\\ c^{+}(\boldsymbol{p})+N&\mbox{ if }\boldsymbol{p}\not\in P^{+}\end{array}\right$.
holds for $N\geq|\boldsymbol{p}|-2$.
(ii) There exists an integer $c^{-}(\boldsymbol{p})$ such that
$f_{N}(\boldsymbol{p})=\left\{\begin{array}[]{ll}c^{-}(\boldsymbol{p})+N&\mbox{ if }\boldsymbol{p}\in P^{-},\\ c^{-}(\boldsymbol{p})-N&\mbox{ if }\boldsymbol{p}\not\in P^{-}\end{array}\right$.
holds for $N\leq-|\boldsymbol{p}|$.
@@
This proposition states that, for example, when $N$ is positive and sufficiently large, the player who takes the last stone is determined by whether the position is in $P^{+}$. Under the constraint that the player taking the last stone must always move to a P-position, the difference in the number of stones taken by both players assuming they take as many stones as possible is given by $c^{+}(\boldsymbol{p})$. For instance, in the position $(5,4,2)$, if $N$ is positive and sufficiently large, the first player will take the last stone. However, to achieve this, the first player must allow the second player to take 7 stones, which is unavoidable (while first player takes 4 stones). Thus, the payoff function is given by $f_{N}(5,4,2)=(4-7)+N=-3+N$, implying that $c^{+}(5,4,2)=-3$.
@@
Proof of Proposition 8. (i) When $\boldsymbol{p}\not\in P^{+}$, the first player has a winning strategy in the normal play Nim (i.e., taking the last stone), ensuring they obtain at least $1+N$ points. On the other hand, if the first player does not take the last stone, the total points obtained by taking all remaining stones is at most $|\boldsymbol{p}|-1$. Thus, if $1+N\geq|\boldsymbol{p}|-1$, which simplifies to $N\geq|\boldsymbol{p}|-2$, the first player will always select to take the last stone. Under the constraint that the first player must always select the next game state from within $P^{+}$, an optimal strategy that maximizes their payoff is determined. The value of $c^{+}(\boldsymbol{p})$ is given by the difference between the number of stones the first player can take and the number the second player can take. If $\boldsymbol{p}\in P^{+}$, then $c^{+}(\boldsymbol{p})$ is determined under the constraint that the second player must always select the next game state from within $P^{+}$.
@@
(ii) In this case, similar to (i), we compare the maximum possible points when a player takes the last stone, $|\boldsymbol{p}|+N$, with the minimum possible points when a player avoids to take it, which is 0. If $|\boldsymbol{p}|+N\leq 0$, the player who wins in the misère play Nim will always select the next game state from within $P^{-}$. Under this constraint, $c^{-}(\boldsymbol{p})$ is determined.  $\Box$
@@
As an example of $c^{+}(\boldsymbol{p})$ and $c^{-}(\boldsymbol{p})$, we compute $c^{+}(3,2,1)$ and $c^{-}(3,2,1)$ in the three-pile Scoring Nim. Starting from $(3,2,1)$, the opponent responds with a move that leads to the next game state in either $P^{+}$ or $P^{-}$ position when $|N|$ is sufficiently large. Specifically, possible moves include $(3,2,1)\to(2,2,1)\to(2,2,0)\in P^{+}\cap P^{-}$, $(3,2,1)\to(1,2,1)\to(1,0,1)\in P^{+}$ or $(1,1,1)\in P^{-}$, among others. By evaluating all possible moves, we find that:
$c^{+}(3,2,1)=3-1+c^{+}(0,1,1)=2,\quad c^{-}(3,2,1)=3-2+c^{-}(0,0,1)=2.$
@@
Similarly, considering all possible outcomes for $(2k+1,2k,1)$ ($k\geq 2$) when $|N|$ is sufficiently large, we find that the optimal sequence of moves consists of the current player first taking 3 stones from the pile with $2k+1$ stones, followed by the opponent taking 1 stone from the pile with $2k$ stones; that is,
$c^{+}(2k+1,2k,1)$ $=$ $3-1+c^{+}(2k-2,2k-1,1),$
$c^{-}(2k+1,2k,1)$ $=$ $3-1+c^{-}(2k-2,2k-1,1).$
This leads to the inductive conclusion that
$c^{+}(2k+1,2k,1)=c^{-}(2k+1,2k,1)=2k\quad(k\geq 1).$ (5.1)
Here the current player effectively implements a strategy in which, through alternating moves, they repeat as many times as possible the situation where they take two more stones than the opponent. This result will be used in the proof of Theorem 11 in the next section.
@@ 6 Changes in the optimal strategy depending on the bonus value
Proof. For (6.1), we apply the recursion formula (3.2) and use induction on the total number of stones to establish the result. We assume here $x>z$.
First, the first player may take all the stones from the pile of size $z$. By (4.3), we have
@@
To prove the reverse inequality, we show that for any move by the first player, the second player can respond so that $f_{N}(x,x,z)$ does not exceed z1+|1+N|.z-1+|1+N|.
  • If the first player moves to $(x,0,z)$, then by (4.4), we have
    $x-f_{N}(x,0,z)=x-\left((x-z)+|1-|1+N||\right)\leq z-(1-|1+N|).$
  • If the first player moves to $(x,x^{\prime},z)$ (where $0<x^{\prime}<x$), the second player responds by moving to $(x^{\prime},x^{\prime},z)$ unless $x^{\prime}=1$ and $z=1$. By the induction hypothesis, it follows that
@@
    If $x^{\prime}=1$ and $z=1$, the second player responds by moving to $(0,1,1)$, and hence
@@
  • If the first player moves to $(x,x,z^{\prime})$ (where $0<z^{\prime}<z$), the second player responds by moving to $(x,x,0)$. Then, by (4.3), we have
@@
Thus, the case $x>z$ has been proved. The cases $x=z$ and $x<z$ follow since the optimal play for the first player is also to move to $(x,x,0)$.  $\Box$
@@
We classify $f_{N}(x,y,1)$ according to cases, where the condition $(x,y,1)\in P^{+}\cap P^{-}$ holds only when $(2k+1,2k,1)$ with $k\geq 1$. We define the function $F_{k}(N)$ for $k\geq 1$ as follows.
Definition 3.
@@
06-66622NN
@@
For example, when $k=3$, $F_{3}(N)$ is depicted in Fig. 3. We record the following properties of $F_{k}(N)$ as a lemma.
Lemma 10.
For $k\geq 1$, the function $F_{k}(N)$ satisfies the following properties:
(i) $F_{k}(N)\leq 2$ for all $N$, and $F_{k}(N)=2$ if and only if $N\in J_{k}$.
(ii) When $N$ is an odd integer with $|N|\leq 2k-1$, then $F_{k}(N)=1$.
(iii) $F_{k}(2k)=F_{k}(-2k)=0$.
(iv) $F_{k}(N)=2k-|N|$ for $|N|\geq 2k-2$.
(v) For all $N$, $\max\{F_{k}(N),-2k+|N|\}\geq 0$ and $\max\{F_{k}(N),-(2k-2)+|N|\}\geq 1$.
@@
Thus, we have fully characterized $f_{N}(x,y,1)$. For example, $f_{N}(8,7,1)$ is an instance of the case where $k=3$ and $x=8$ in part (iii) of the theorem above. It is given by $\max\{F_{3}(N),-4+|N|\}$, which is represented by the solid line in Fig. 4.
$\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot$$\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot$$\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot$$\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot$$\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot$$\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot\cdot$06-64-4446622NN
Figure 4: $f_{N}(8,7,1)=\max\{F_{3}(N),-4+|N|\}$ (solid line)
One key observation concerns the number of breakpoints in $f_{N}(x,y,1)$, namely, points where the slope changes sign. We find that $f_{N}(2k+1,2k,1)$ has 4k34k-3 breakpoints, indicating that the number of breakpoints can grow arbitrarily large as the number of stones increases.
Proof of Theorem 11. For $k=1$, (i), (ii), and (iii) can be verified directly. For $k\geq 2$, we proceed by induction, assuming that (i), (ii), and (iii) hold for all $k<m$, and proving that (i) holds for $k=m$. Under this assumption, the following facts (a), (b), and (c) hold for $1\leq k<m$.
(a) For all $N$, we have $f_{N}(2k+1,2k,1)\leq 2$, with equality holding for all $N$ in $J_{k}$. In particular, if $N$ is an odd integer, then $f_{N}(2k+1,2k,1)\leq 1$. Moreover, if $|N|\geq 2k-2$, then $f_{N}(2k+1,2k,1)=2k-|N|$. Consequently, $f_{N}(2k+1,2k,1)\leq-1$ if $|N|\geq 2k+1$ and $f_{N}(2k+1,2k,1)\leq-2$ if $|N|\geq 2k+2$.
(b) For $x\geq 2k+2$, it follows that $f_{N}(x,2k,1)\geq x-2k-1$ for all $N$, with equality holding only when $N=\pm 2k$. Moreover, when $N=\pm(2k+1)$, we have $f_{N}(x,2k,1)=x-2k$.
(c) For $x\geq 2k+2$, it follows that $f_{N}(x,2k+1,1)\geq x-2k-1$ for all $N$.
These facts follow easily from Lemma 10 and will be used repeatedly in this proof. Notably, in (b), equality holds for only two values of $N$. Specifically, $f_{N}(6,4,1)$ is an instance of the case where $k=2$ and $x=6$ in Theorem 11 (ii), and as illustrated in Fig. 5, it attains its minimum value at $N=\pm 4$.
04-44433NN
@@
As stated in Section 3, Proposition 1(ii) implies that $f_{N}(p)$ is a continuous piecewise linear function of $N$ with slope $\pm 1$. Hence, to establish (i), it suffices to consider integer values of $N$. We outline the proof strategy as follows:
  • By showing that $f_{N}(2m+1,2m,1)\geq 2$ when $N$ is an even integer in the range $|N|\leq 2m-2$, it follows that $f_{N}(2m+1,2m,1)\geq F_{m}(N)$ for all $N$.
  • By showing that $f_{N}(2m+1,2m,1)\leq 1$ when $N$ is an odd integer in the range $|N|\leq 2m-3$, and using (5.1), it follows that $f_{N}(2m+1,2m,1)\leq F_{m}(N)$ for all $N$.
We will explain in detail in the proof below that it is sufficient to establish these facts.
Case 1: $N=\pm 2l$ with $0\leq l<m$. The first player has the option to move from the position $(2m+1,2m,1)$ to $(2l,2m,1)$. By the induction hypothesis—namely, by applying Proposition 3 when $l=0$, and using (b) when $1\leq l<m$—we have $f_{\pm 2l}(2l,2m,1)=2m-2l-1$. Hence we obtain